OpenStudy (nuccioreggie):

Which expression is equivalent to two-thirds divided by five-sixths? five-sixths times two-thirds two-thirds times five-sixths three-halves times six-fifths two-thirds times six-fifths

OpenStudy (nuccioreggie):

@chosenmatt

OpenStudy (nuccioreggie):

@HeatDude

OpenStudy (gorv):

\[\frac{ 2 }{ 3 } \div \frac{ 5 }{ 6 }\]

OpenStudy (gorv):

\[=\frac{ 2 }{ 3 }*\frac{ 6 }{ 5 }=\frac{ 2*6 }{ 3*5 }\]

OpenStudy (nuccioreggie):

@mathmath333

OpenStudy (gorv):

i think u can do it now

OpenStudy (nuccioreggie):

@StudyGurl14

OpenStudy (nuccioreggie):

@gorv so its B right

OpenStudy (studygurl14):

No. Using the information gorv gave, it would be D.

OpenStudy (gorv):

Agree WITH @StudyGurl14